OpenStudy (anonymous):

If an animal is h feet high at the shoulder, then the number N of steps per second that the animal takes while walking can be estimated by N(h) =1.6h^-1/2. Estimate the stepping frequency of a dog that is 2.9 feet high at the shoulders.

OpenStudy (shane_b):

They give you the formula...just plug in the dog's height:\[N(h)=1.6h^{-\frac{1}{2}}\]\[h=2.9ft\]\[N(h)=1.6(2.9ft)^{-\frac{1}{2}}\]